# Beyond the Silence: How Men Navigate Infertility Through Digital Communities and Data Sharing

### v1

@@ 5.1.3. Bridge-builders who normalize norms and transfer know-how
When answering RQ3, we found cross-posters resemble collective-oriented brokers: they write generalizable, guidance-seeking content, lean on “we” pronouns, and often activate when engagement is limited locally—complementing prior observations of cross-community knowledge brokering and empathy building especially when managing trauma (Ammari et al., 2025; Randazzo and Ammari, 2025). This finding also complicates the study of norm creation at different community levels (i.e., The Reddit platform $\rightarrow$ several related subreddits $\rightarrow$ one subreddit $\rightarrow$ micro-communities within the subrreddit) (Chandrasekharan et al., 2018) as it shows how cross-posters can engage in norm-setting in different communities relying on their deep knowledge of community norms in both spaces.
@@ 5.2.3. Visible role and expertise tags—with optional professional verification
Building on our role findings in §4.2.1 and prior work on badges, awards, and verified participation (Burtch et al., 2022; Trujillo, 2022; Bornfeld and Rafaeli, 2017) in §2.3.2, we recommend a multilayered tagging and verification system to enhance visibility and trust within the community.
@@
Complementing these role tags, members could voluntarily adopt topic or skill tags (e.g., insurance navigation, TESE recovery) to make their expertise legible to seekers. In parallel, professionally verified accounts for clinicians and counselors would strengthen credibility, following verification approaches such as license checks and scoped disclaimers proposed in (Ammari et al., 2025). Importantly, these accounts must include clear guardrails to prevent their presence from being misconstrued as a substitute for clinical care.
@@ A.1.1. Appendix: Hyperparameter Selection
    Minimum Topic Size: The min_topic_size parameter controls the smallest allowable topic size in the clustering process (Grootendorst, 2020). We explored values of {20, 30, 40, 50}:

### v2

@@ 5.1.3. Bridge-builders who normalize norms and transfer know-how
When answering RQ3, we found cross-posters resemble collective-oriented brokers: they write generalizable, guidance-seeking content, lean on “we” pronouns, and often activate when engagement is limited locally—complementing prior observations of cross-community knowledge brokering and empathy building especially when managing trauma (Ammari et al., 2025; Randazzo and Ammari, 2025). This finding also complicates the study of norm creation at different community levels (i.e., The Reddit platform $\rightarrow$ several related subreddits $\rightarrow$ one subreddit $\rightarrow$ micro-communities within the subrreddit) (Chandrasekharan et al., 2018) as it shows how cross-posters can engage in norm-setting in different communities relying on their deep knowledge of community norms in both spaces. Additionally, as we showed in §4.1.3, the norms differences between r/maleinfertility and r/infertility are contested, which raises questions about how these differences are understood over time and how cross-posters affect the evolution of both communities.
@@ 5.2.3. Visible role and expertise tags—with optional professional verification
Building on our role findings in §4.4 and prior work on badges, awards, and verified participation (Burtch et al., 2022; Trujillo, 2022; Bornfeld and Rafaeli, 2017) in §2.3.2, we recommend a multilayered tagging and verification system to enhance visibility and trust within the community.
@@
Complementing these role tags, members could voluntarily adopt topic or skill tags (e.g., insurance navigation, azoospermia recovery) to make their expertise legible to seekers. In parallel, professionally verified accounts for clinicians and counselors would strengthen credibility, following verification approaches such as license checks and scoped disclaimers proposed in (Ammari et al., 2025). Importantly, these accounts must include clear guardrails to prevent their presence from being misconstrued as a substitute for clinical care.
@@ A.1.1. Appendix: Hyperparameter Selection
    Minimum Topic Size: The min_topic_size parameter controls the smallest allowable topic size in the clustering process (Grootendorst, 2022). We explored values of {20, 30, 40, 50}: